OpenStudy (anonymous):

If $13,300 is invested at 3.7% interest compounded semi-annually, how much will the investment be worth in 18 years? A. $25,730.54 B. $25,808.47 C. $25,861.09 D. $25,578.16

OpenStudy (anonymous):

\[I = P(1 + \frac rn)^{nt}\]

OpenStudy (anonymous):

Where: P = Principal r = rate in decimals n = number of times compounded a year t = number of years

OpenStudy (anonymous):

so wat may be the answer

OpenStudy (anonymous):

You're gonna have to plug the values into the formula.

OpenStudy (anonymous):

I can't just give you the answers.

OpenStudy (anonymous):

I = 13300(1 + 0.037/2)^36 Can you solve it now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ican't resolve it i dont have parenthesis

OpenStudy (anonymous):

Just do what's in the parentheses first. I = 13300(1 + 0.037/2)^36 I = 13300(1 + 0.0185)^36 I = 13300(1.0185)^36 Now just plug in 1.0185^36 and then multiply that by 13300

OpenStudy (anonymous):

omg is too complicated

OpenStudy (anonymous):

It's not complicated at all. It just looks complicated.

OpenStudy (anonymous):

idk man or girl about this problem

OpenStudy (anonymous):

Do you have a scientific calculator?

OpenStudy (anonymous):

no i dont

OpenStudy (anonymous):

thats why

OpenStudy (anonymous):

Do you have a regular calculator?

OpenStudy (anonymous):

yeah in the computer

OpenStudy (anonymous):

Just multiply 1.9346268 by 13300 using the calculator on your computer.

OpenStudy (anonymous):

Post your answer on here.

OpenStudy (anonymous):

is a

OpenStudy (anonymous):

25730.53

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

alright thank u :) i will looking for u if i need ur help :)

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

http://tinypic.com/view.php?pic=24xhqmt&s=6 this one plz

OpenStudy (anonymous):

I believe that it's C.

OpenStudy (anonymous):

really ah ok

OpenStudy (anonymous):

Rudolph's average balance checking account pays simple interest of 8.4% annually, and he made $6.93 in interest last month. What was Rudolph's average balance last month? A. $9900.00 B. $99.00 C. $9.90 D. $990.00

OpenStudy (anonymous):

whbu this one

OpenStudy (anonymous):

I'm not completely sure, but I think that it's D. I gotta go, so ask someone to double check my work.

OpenStudy (anonymous):

Pablo deposited $57 into a savings account for which interest is compounded quarterly. According to the Rule of 72, what interest rate will cause his money to double in approximately 44 years? this one A. 1.3% B. 0.8% C. 0.6% D. 1.6%